# On the Fourier Coefficients of Critical Gaussian Multiplicative Chaos

### v1

@@ 2.2. Brownian estimates
We record the following estimates for the probability that a Brownian motion stays under a barrier starting at time $e$ (in light of (2.9)). Such events will arise from restricting to samples $\omega$ for which the event in Corollary 2.5 holds.
@@
Let $B_{t}$ be a standard linear Brownian motion and $a>0$. Then for any $t>e$,
(2.10) $\mathbb{P}\Big(\sup_{s\in[e,t]}B_{s}\leq a\Big)\ll\frac{(a+1)}{\sqrt{t}+1}.$
@@
(2.11) $\mathbb{P}_{(0,a)}^{(t,b)}\Big(B_{s}\geq 0,\forall s\in[e,t]\Big)\ll\frac{(a+1)(b+1)}{t+1}.$
@@ 3. Proof of the main theorem
[…]
We want to show that $\{(\log n)\cdot|c_{n}|\}_{n\geq 1}$ is tight, which we do by proving an upper bound of the correct order for the expectation of
(3.2) $|c_{n,t}|^{2}=\iint_{[0,1]^{2}}e^{in(\theta_{2}-\theta_{1})}\mu_{t}(\mathrm{d}\theta_{1})\mu_{t}(\mathrm{d}\theta_{2}).$
on a good event $\mathcal{E}_{n,t}^{(A)}$. This is the content of Proposition 3.1 below. As a byproduct of its proof, we identify the set of pairs $(\theta_{1},\theta_{2})\in[0,1]^{2}$ over which the integral in (3.2) yields a leading order contribution, and those for which the contribution is negligible in the limit (see Remark 4.4).
@@
$\mathcal{E}^{(A)}_{n,t}:=\big\{\forall\theta\in[0,1],\,\theta\in\mathcal{G}_{n,t}^{(A)}\big\},$
@@
(3.3) $\mathcal{G}_{n,t}^{(A)}=\left\{\theta\in[0,1]:\,\,\begin{aligned} &X_{s}^{(r_{n})}(\theta)\leq U(s)+A,\,\forall s\in[0,t-r_{n}]\\ &X_{r_{n}}(\theta)\leq m(r_{n})+A\end{aligned}\right\}$
and $m(t),U(t)$ are defined in Equations (2.8), (2.9) respectively. Then there exists an integer $N>2$ such that for any $n\geq N$,
$\limsup_{t\to\infty}\mathbb{E}\Big[|c_{n,t}|^{2};{\mathcal{E}_{n,t}^{(A)}}\Big]\ll_{A}(\log n)^{-2},$
@@
The choice of good event $\mathcal{E}_{n,t}^{(A)}$ will be explained in Section 4. For the time being, we show how Theorem 1.1 follows straightforwardly from this proposition.
@@
Noting that $|c_{n}|\leq\mu([0,1])$ is almost surely finite, it suffices to prove that $\{|c_{n}|\log n\}_{n\geq N}$ is tight where $N$ is the constant from Proposition 3.1.
Fix $\epsilon>0$ and let $K>0$ be arbitrary for now. Then by (3.1) and the Portmanteau theorem, we have that
$\mathbb{P}\Big(|c_{n}|(\log n)>K\Big)$ $\leq\limsup_{t\to\infty}\mathbb{P}\Big(|c_{n,t}|(\log n)>K\Big)$
for any $n>0$. Furthermore, by definition of $\mathcal{E}_{n,t}^{(A)}$, the pre-limit quantity on the right-hand side is bounded above by
$\mathbb{P}\Big(\mathbf{1}_{\mathcal{E}_{n,t}^{(A)}}$ $|c_{n,t}|(\log n)>K\Big)+\mathbb{P}\Big(\exists\theta\in[0,1]:X_{r_{n}}(\theta)>m(r_{n})+A\Big)$
(3.4) $+\mathbb{P}\Big(\exists s\in[0,\infty),\exists\theta\in[0,1]:X_{s}^{(r_{n})}(\theta)>U(s)+A\Big),$
for any $A>0$. We can then make the second and third terms in (3) arbitrarily small in a way that is uniform in $n$, by taking $A$ large enough. Indeed, for any $n$, $(X_{t}^{(r_{n})})_{t\geq 0}$ (defined in (2.7)) is a sequence of cutoff approximations to a $\star$–scale invariant field with variance $t$ at every point. Corollary 2.5 and Proposition 2.3 thus imply that
(3.5) $\sup_{n>2}\mathbb{P}\Big(\exists s\in[0,\infty),\exists\theta\in[0,1]:X_{s}^{(r_{n})}(\theta)>U(s)+A\Big)<\varepsilon/3,$
(3.6) $\sup_{n>2}\mathbb{P}\Big(\exists\theta\in[0,1]:X_{r_{n}}(\theta)>m(r_{n})+A\Big)<\varepsilon/3,$
for sufficiently large $A$. For such an $A$, it follows that
$\sup_{n\geq N}\mathbb{P}\Big(|c_{n}|(\log n)>K\Big)$ $\leq\sup_{n\geq N}\limsup_{t\to\infty}\mathbb{P}\Big(\mathbf{1}_{\mathcal{E}_{n,t}^{(A)}}|c_{n,t}|(\log n)>K\Big)+2\varepsilon/3.$
Using Chebyshev’s inequality and Proposition 3.1, we can then write
$\sup_{n\geq N}\limsup_{t\to\infty}\mathbb{P}\Big(\mathbf{1}_{\mathcal{E}_{n,t}^{(A)}}|c_{n,t}|(\log n)>K\Big)$
$\leq\sup_{n\geq N}\frac{(\log n)^{2}}{K^{2}}\limsup_{t\to\infty}{\mathbb{E}\big[|c_{n,t}|^{2};\mathcal{E}_{n,t}^{(A)}\big]}\ll_{A}\frac{1}{K^{2}}.$
Tightness follows by picking $K=K(A)$ to make the left-hand side $\varepsilon/3$. ∎
What’s left is to prove Proposition 3.1, which is the main difficulty here.
@@ 4.1. Initial reductions
To bound $|c_{n,t}|\mathbf{1}_{\mathcal{E}_{n,t}^{(A)}}$, we first want to extract a local condition (meaning a condition on each pair of points $\theta_{1},\theta_{2}$) in (3.2) from the global event $\mathcal{E}_{n,t}^{(A)}$, noting that one cannot directly upper bound the integrand due to the presence of the oscillating factor $e^{in\Delta}$. We instead note the following pointwise inequality on the sample space
(4.1) $\mathbf{1}_{\mathcal{E}_{n,t}^{(A)}}|c_{n,t}|^{2}$ $=\mathbf{1}_{\mathcal{E}_{n,t}^{(A)}}\iint_{[0,1]^{2}}e^{in(\theta_{2}-\theta_{1})}\mathbf{1}\Big\{\theta_{1},\theta_{2}\in\mathcal{G}_{n,t}^{(A)}\Big\}\mu_{t}(\mathrm{d}\theta_{1})\mu_{t}(\mathrm{d}\theta_{2})$
$\leq\iint_{[0,1]^{2}}e^{in(\theta_{2}-\theta_{1})}\mathbf{1}\Big\{\theta_{1},\theta_{2}\in\mathcal{G}_{n,t}^{(A)}\Big\}\mu_{t}(\mathrm{d}\theta_{1})\mu_{t}(\mathrm{d}\theta_{2}).$
@@
$\tilde{\mu}_{t}(\mathrm{d}\theta):=\mathbf{1}\Big\{\theta\in\mathcal{G}_{n,t}^{(A)}\Big\}\mu_{t}(\mathrm{d}\theta).$
@@
(4.4) $\mathcal{C}_{n,t}(\#):=\iint_{(\mathrm{\#})}e^{in(\theta_{2}-\theta_{1})}\mathbf{1}\Big\{\theta_{1},\theta_{2}\in\mathcal{G}_{n,t}^{(A)}\Big\}\mu_{t}(\mathrm{d}\theta_{1})\mu_{t}(\mathrm{d}\theta_{2})$
@@
4.2. The oscillatory range
We first study the contribution from $(\mathrm{I)}$ where $e^{in\Delta}$ oscillates very rapidly.

### v2

@@ 2.2. Brownian estimates
We record the following estimates for the probability that a Brownian motion stays under a barrier starting at time $e$ (in light of (2.6)). Such events will arise from restricting to samples for which the event in Proposition 2.2 holds.
@@
Let $B_{t}$ be a standard linear Brownian motion. Then there exists a constant $C>0$ such that for any $t>e$, $a>0$,
(2.7) $\mathbb{P}\Big(\sup_{s\in[e,t]}B_{s}\leq a\Big)\leq C\frac{(a+1)}{\sqrt{t}+1}.$
@@
(2.8) $\mathbb{P}_{(0,a)}^{(t,b)}\Big(B_{s}\geq 0,\forall s\in[e,t]\Big)\leq C\frac{(a+1)(b+1)}{t+1}.$
@@ 3. Proof of the main theorem
(3.1) $c_{n,t}=\int_{0}^{1}e^{in\theta}\mu_{t}(\mathrm{d}\theta)\overunderset{\mathbb{P}}{t\to\infty}{\longrightarrow}c_{n}=\int_{0}^{1}e^{in\theta}\mu(\mathrm{d}\theta).$
We want to show that $\{(\log n)^{\alpha}|c_{n}|\}_{n\geq 1}$ converges to zero in probability for any $\alpha<1/4$, which we do by proving an upper bound of the correct order for the expectation of
(3.2) $|c_{n,t}|^{2}=\iint_{[0,1]^{2}}e^{in(\theta_{2}-\theta_{1})}\mu_{t}(\mathrm{d}\theta_{1})\mu_{t}(\mathrm{d}\theta_{2})$
on a good event ${E}_{n,t}{(A)}$. This is the content of Proposition 3.1 below. As a byproduct of its proof, we identify the set of pairs $(\theta_{1},\theta_{2})\in[0,1]^{2}$ over which the integral in (3.2) yields a leading order contribution, and those for which the contribution is negligible in the limit (see Remark 4.4).
@@
${E}_{n,t}{(A)}:=\big\{\forall\theta\in[0,1],\,\theta\in G_{n,t}{(A)}\big\},$
@@
(3.3) $G_{n,t}{(A)}=\left\{\theta\in[0,1]:\,\,X_{s}(\theta)\leq U(s)+A,\,\forall s\in[r_{n},t]\right\}$
and $U(t)$ is defined in Equation (2.6). Then for any $\varepsilon>0$, there exists an integer $N>2$ such that for any $n\geq N$,
$\limsup_{t\to\infty}\mathbb{E}\Big[|c_{n,t}|^{2};{{E}_{n,t}{(A)}}\Big]\ll_{A}(\log n)^{-1/2+\varepsilon},$
@@
The choice of good event ${E}_{n,t}{(A)}$ will be explained in Section 4. For the time being, we show how Theorem 1.1 follows straightforwardly from this proposition.
@@
Fix $\varepsilon>0$ and $\alpha<1/4$. By (3.1) and the Portmanteau theorem,
$\mathbb{P}\Big(|c_{n}|({\log n})^{\alpha}>\varepsilon\Big)$ $\leq\liminf_{t\to\infty}\mathbb{P}\Big(|c_{n,t}|({\log n})^{\alpha}>\varepsilon\Big)$
for any $n>0$. By definition of ${E}_{n,t}{(A)}$, the pre-limit quantity on the right-hand side is bounded above by
(3.4) $\mathbb{P}\Big(\mathbf{1}_{{E}_{n,t}{(A)}}$ $|c_{n,t}|({\log n})^{\alpha}>\varepsilon\Big)+\mathbb{P}\Big(\exists s\in[0,\infty),\exists\theta\in[0,1]:X_{s}(\theta)>U(s)+A\Big),$
for any $A>0$. By Proposition 2.2, we can make the second term in (3.4) arbitrarily small in a way that is uniform in $n$. We have that
(3.5) $\sup_{n>2}\mathbb{P}\Big(\exists s\in[0,\infty),\exists\theta\in[0,1]:X_{s}(\theta)>U(s)+A\Big)<\varepsilon/2,$
for sufficiently large $A$, and it follows that for such an $A$ and any $n>0$,
$\mathbb{P}\Big(|c_{n}|({\log n})^{\alpha}>\varepsilon\Big)$ $\leq\limsup_{t\to\infty}\mathbb{P}\Big(\mathbf{1}_{{E}_{n,t}{(A)}}|c_{n,t}|({\log n})^{\alpha}>\varepsilon\Big)+\varepsilon/2.$
Using Chebyshev’s inequality, we can then write
$\limsup_{t\to\infty}\mathbb{P}\Big(\mathbf{1}_{{E}_{n,t}{(A)}}|c_{n,t}|({\log n})^{\alpha}>\varepsilon\Big)\leq\frac{(\log n)^{2\alpha}}{\varepsilon^{2}}\limsup_{t\to\infty}{\mathbb{E}\big[|c_{n,t}|^{2};{E}_{n,t}{(A)}\big]},$
and the claim follows by Proposition 3.1 upon taking $n\to\infty$. ∎
What’s left is to prove Proposition 3.1, which is the main task here.
@@ 4.1. Initial reductions
To bound $|c_{n,t}|\mathbf{1}_{{E}_{n,t}{(A)}}$, we first want to extract a local condition (meaning a condition on each pair of points $\theta_{1},\theta_{2}$) in (3.2) from the global event ${E}_{n,t}{(A)}$, noting that one cannot directly upper bound the integrand due to the presence of the oscillating factor $e^{in(\theta_{2}-\theta_{1})}$. We instead note the following pointwise inequality on the sample space
(4.1) $\mathbf{1}_{{E}_{n,t}{(A)}}|c_{n,t}|^{2}$ $=\mathbf{1}_{{E}_{n,t}{(A)}}\iint_{[0,1]^{2}}e^{in(\theta_{2}-\theta_{1})}\mathbf{1}\Big\{\theta_{1},\theta_{2}\in G_{n,t}{(A)}\Big\}\mu_{t}(\mathrm{d}\theta_{1})\mu_{t}(\mathrm{d}\theta_{2})$
$\leq\iint_{[0,1]^{2}}e^{in(\theta_{2}-\theta_{1})}\mathbf{1}\Big\{\theta_{1},\theta_{2}\in G_{n,t}{(A)}\Big\}\mu_{t}(\mathrm{d}\theta_{1})\mu_{t}(\mathrm{d}\theta_{2}).$
@@
$\tilde{\mu}_{t}(\mathrm{d}\theta):=\mathbf{1}\Big\{\theta\in G_{n,t}{(A)}\Big\}\mu_{t}(\mathrm{d}\theta).$
@@
(4.4) $\mathcal{C}_{n,t}(\#):=\iint_{(\mathrm{\#})}e^{in(\theta_{2}-\theta_{1})}\mathbf{1}\Big\{\theta_{1},\theta_{2}\in G_{n,t}{(A)}\Big\}\mu_{t}(\mathrm{d}\theta_{1})\mu_{t}(\mathrm{d}\theta_{2})$
@@
4.2. The oscillatory range ($|\theta_{2}-\theta_{1}|\geq\Delta_{n}$)
We first study the contribution from $(\mathrm{I)}$ where $e^{in(\theta_{2}-\theta_{1})}$ oscillates very rapidly.